OpenStudy (anonymous):

how do you do cos2x=0, the answer says 3pi/4

OpenStudy (jhannybean):

Depends if you're integrating or derivating

OpenStudy (anonymous):

oh, sorry, derivating.

OpenStudy (ddcamp):

cos(2x) = 0 2x = arcos(0) 2x = 3pi / 2 x = 3pi/4

OpenStudy (jhannybean):

\[\frac{ d }{ dx }\cos2x\]

OpenStudy (anonymous):

what is arcos?

OpenStudy (jhannybean):

Yeah...DD really knows his unit circle.

OpenStudy (jhannybean):

arc cos is the inverse cosine used to find the angle of measurement

OpenStudy (anonymous):

cos equals o at pi/2 and 3pi/2?

OpenStudy (jhannybean):

arc cos = \[\cos ^{-1}(\theta)\]

OpenStudy (ddcamp):

arccos is the inverse cosine function: arccos(0) is the values of theta where cos(theta) = 0

OpenStudy (anonymous):

yes, but cos equals 0 at pi/2 and 3pi/2 so how do you know which one to use?

OpenStudy (anonymous):

In general, when : \( cos(x) = 0\) Then: \[x = (2n+1) \frac{\pi}{2}\] where n belongs to set of Integers..

OpenStudy (ddcamp):

It depends. Usually the question will include a range of possible theta values, either from 0 to pi, or 0 to 2pi, or something. If they don't give you the range, either value should work, even if it's not the answer in the book.

OpenStudy (jhannybean):

You'd know which one to use depending on the angle of measurement you're looking for.

OpenStudy (jhannybean):

Or which radians you're given and are working towards finding the angle measurement

OpenStudy (anonymous):

what if i was finding a minimum value from an equation? would i just list both?

OpenStudy (jhannybean):

Im not sure what you mean by minimum value?

OpenStudy (anonymous):

i don't understand how DDCamp got from.. 2x = arcos(0) to 2x = 3pi / 2 does arcos(0) just mean where cos is equal to 0? we haven't covered ar? what does it mean?

OpenStudy (jhannybean):

I'mout, @genius12 , @waterineyes or @DDCamp will helpyou out with your problems! :D Gooooood luck!!!

OpenStudy (ddcamp):

Yes, arccos(0) is where cos(theta)=0

OpenStudy (anonymous):

Who took your wicket @Jhannybean or you are Retired Hurt??? :P

OpenStudy (jhannybean):

Yes, DDcamp should explain why he chose to use \[\frac{ 3\pi }{ 2 }\] instead of \[\frac{ \pi }{ 2 }\] to solve for x

OpenStudy (jhannybean):

lol @waterineyes , I have to go study for a final!!! Engllish...=_=

OpenStudy (anonymous):

From what it looks like, I believe you're solving for x. And I'm also assuming that your solutions are to exist in the interval [0, 2pi]. If the aforementioned conditions hold, then this is how you would solve for x:\[\bf \cos(2x)=0\]The only places where Cosine is 0 in the given interval is at pi/2 and 3pi/2, which means:\[\bf \implies 2x=\frac{\pi}{2} \ or \ 2x=\frac{3 \pi}{2} \implies x=\frac{\pi}{4},\frac{3 \pi}{4}\] Do you understand?

OpenStudy (anonymous):

Hey use the above one that I gave, put n = 1 there.. You can put 0, 1, 2 or -1, -2 all but according to your answer n = 1 there..

OpenStudy (jhannybean):

@genius12 found the write explanation for this problem :)

OpenStudy (anonymous):

\[\cos(2x) = 0\] \[\implies 2x = (2n+1)\frac{\pi}{2}\] Put n= 1 here : \[2x = \frac{3\pi}{2}\] \[x = \frac{3\pi}{4}\]

OpenStudy (anonymous):

okay, i understand, thankyou all!

OpenStudy (anonymous):

how do you solve y=sin(2x+pi/4) to find the tangent and normal when x=0